FOURIER TRANSFORM NEAR INFRARED EMISSION SPECTROSCOPY OF GAS-PHASE Y bO

TODD C. MELVILLE, ^{a}Department of Chemistry, Dalhousie University, Halifax, Nova Scotia, B3H 4J3, Canada; \underline{IOULI GORDON}, KEITH A. TERESZCHUK, PETER F. BERNATH, Guelph-Waterloo Centre for Graduate Work in Chemistry and Biochemistry, University of Waterloo, Waterloo, Ontario, N2L 3G1 Canada; JOHN A. COXON, Department of Chemistry, Dalhousie University, Halifax, Nova Scotia, B3H 4J3, Canada.

The emission spectrum of gas-phase YbO has been investigated using a Fourier transform spectrometer. Chemiluminescence was observed from excited YbO molecules produced in a Broida-type oven by the reaction of ytterbium metal vapour with N_{2}O. A total of 8 red-degraded bands in the range 9800 - 11300 cm^{−1} were recorded at a resolution of 0.04 cm^{−1}. Because of the multiple isotopomers present in the spectra, only 3 bands were rotationally analyzed. Perturbations were identified in two of these bands and all 3 transitions were found to terminate at the X ^{1} Σ ^{+} ground electronic state. The electronic configurations that give rise to the observed states are investigated and molecular parameters for all of the analyzed bands will be presented.

aPresent address:Department of Chemistry, University of British Columbia, Vancouver, British Columbia, V6T 1Z1,Canada